A wire is stretched by 8mm when a load of 60N is applied. What will be the extension of a wire of the same material having four times the cross sectional area and twice the original length when one eighth of the load is applied
1
Expert's answer
2020-01-27T09:32:12-0500

We have:


"\frac{Fl}{Ax}=\frac{F(2l)}{8(4A)x'}"

"x'=\frac{1}{16}x=\frac{1}{16}8=0.5\ mm"